OpenStudy (anonymous):

Simplify. 1+[2x-3(1.8/0.2)] A.2x-26 B.2x+28 C.3x-27 D.-7x+1

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@CausticSyndicalist

OpenStudy (anonymous):

d

OpenStudy (causticsyndicalist):

I think it would be better to explain it. Anyways, does it lay out like this? \[1+2x-3\left(\begin{matrix}1.8 \\ 0.2\end{matrix}\right)\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[1+[2x-3(1.8 \div 0.2)]\]

OpenStudy (causticsyndicalist):

Ohhh okay. and since the question wants it to be simplified, the first step would be to simplify \[\frac{ 1.8 }{ 0.2 }\] to 9

OpenStudy (anonymous):

okay then what?

OpenStudy (causticsyndicalist):

Then you simply 3 x 9 to 27.

OpenStudy (anonymous):

okay then we add one?

OpenStudy (causticsyndicalist):

I think we just simply 2x - 26 from here. Not sure about this step. Maybe you could try this and then maybe try with with adding the one. See which one comes up with one of the answers in the multiple choice list.